OpenStudy (henrietepurina):

Help for a medal! Calculate the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of 1/2.

OpenStudy (acxbox22):

y+3=1/2(x-6)

OpenStudy (henrietepurina):

thanks

OpenStudy (anonymous):

it is essential that these formulas are known slope formula y - y1 m = ----- x - x1 point slope formula y - y1 = m(x - x1) slope intercept y=mx+b